Lydia makes a calculation and gets an answer of 22.5. On the last step, she multiplies by 0.3 when she should have divided by 0.3. What should the correct answer be?
Answer by htmentor(1343)

Let x = the number she had prior to the last operation
Then 0.3x = 22.5 -> x = 22.5/0.3 = 75
She should have divided x by 0.3, so the correct answer is:
75/0.3 = 250